# AeroGPT: Leveraging Large-Scale Audio Model for Aero-Engine Bearing Fault Diagnosis

### v1

@@ Abstract
Aerospace engines, as critical components in aviation and aerospace industries, require continuous and accurate fault diagnosis to ensure operational safety and prevent catastrophic failures. While deep learning techniques have been extensively studied in this context, they output logits or confidence scores, necessitating post-processing to derive actionable insights. Furthermore, the potential of large-scale audio models in this domain remains largely untapped. To address these limitations, this paper proposes AeroGPT, a novel framework that transfers knowledge from general audio domain to aero-engine bearing fault diagnosis. AeroGPT is a framework based on large-scale audio model that incorporates Vibration Signal Alignment (VSA) to adapt general audio knowledge to domain-specific vibration patterns, and combines Generative Fault Classification (GFC) to directly output interpretable fault labels. This approach eliminates the need for post-processing of fault labels, supports interactive, interpretable, and actionable fault diagnosis, thereby greatly enhancing industrial applicability. Through comprehensive experimental validation on two aero-engine bearing datasets, AeroGPT achieved exceptional performance with 98.94% accuracy on the DIRG dataset and perfect 100% classification on the HIT bearing dataset, surpassing traditional deep learning approaches. Additional Qualitative analysis validates the effectiveness of our approach and highlights the potential of large-scale models to revolutionize fault diagnosis.
Index Terms:
Aerospace Engine, Deep Learning, Large Language Model, Fault Diagnosis
@@ I Introduction
Aerospace engines serve as the cornerstone of modern aviation and aerospace industries, powering everything from commercial aircraft to space exploration vehicles [1, 2]. Their reliable operation is paramount to aviation safety, mission success, and the prevention of catastrophic failures that could result in significant human and economic losses. Within these sophisticated propulsion systems, aero-engine bearings are particularly critical components, operating under extreme conditions of high speeds, temperature variations, and significant mechanical stress [3, 4, 5]. The failure of these bearings can lead to severe engine malfunctions, making them a primary focus for condition monitoring and maintenance protocols [6, 2]. Fault diagnosis of aero-engine bearings has consequently emerged as an essential discipline, offering methodologies to detect, identify, and predict potential failures before they manifest as catastrophic events. Early and accurate bearing fault detection can significantly extend engine lifespan, optimize maintenance schedules, and ensure operational safety [7]. Traditionally, bearing fault diagnosis approaches rely on a two-step process of feature extraction followed by classification [8]. For instance, Medina et al. [9] proposed an approach using Poincare plots for feature extraction, with subsequent classification performed by a multi-class Support Vector Machine (SVM). However, such approaches are limited by their dependence on domain-specific feature engineering and conventional machine learning classifiers, which may fail to fully capture the complex dynamics of aero-engine bearing systems.
In recent years, data-driven methods, particularly deep learning approaches, have revolutionized fault diagnosis by enabling end-to-end learning directly from raw vibration signals [10, 11, 12]. These methods automatically extract hierarchical features, eliminating the need for manual feature engineering that traditional approaches require. Convolutional Neural Networks (CNNs) have been widely adopted for this purpose, as they excel at capturing local patterns and spatial hierarchies in data. Wang et al. [13] proposed a one-dimensional CNN with attention mechanism tailored for bearing fault enhancement and classification, establishing a new state-of-the-art performance on the wheelset bearing dataset. Apart from CNNs, Recurrent Neural Networks, especially Long Short-Term Memory (LSTM) networks, have also gained traction in vibration signal analysis. LSTMs are adept at capturing temporal dependencies in sequential data, making them suitable for modeling the time-varying nature of vibration signals. Song et al. [14] developed an CNN-BiLSTM network that combines CNN’s spatial feature extraction capabilities with BiLSTM’s temporal modeling, demonstrating robust performance even with limited samples. Transformer-based architectures [15], leveraging self-attention mechanisms to capture long-range dependencies and contextual information, have also emerged as powerful tools for vibration signal analysis. For instance, Li et al. [16] proposed Dconformer, a novel CNN-Transformer network that combines joint-learning denoising with a multi-branch cross-cascaded architecture to extract both local and global features from noisy vibration signals. Xiang et al. [17] introduced a frequency channel-attention based Vision Transformer method, which integrates frequency domain characteristics with self-attention mechanisms to enhance model interpretability while maintaining computational efficiency through a shift linear layer design. Alongside architectural innovations, significant research has focused on learning strategies to enhance model adaptability and reduce data dependency. Prominent examples include transfer learning, which adapts models pre-trained on large source domains [18, 19]; meta-learning [20] and the closely related field of few-shot learning [21, 22], which aim to train models capable of recognizing new fault types from very limited data; and unsupervised learning [23], which leverages large amounts of unlabeled data.
@@
However, the data-driven methods mentioned above typically output logits or confidence scores, necessitating post-processing to derive actionable insights, whereas industrial applications require efficient and interpretable solutions. Recently, the emrgency and development of large language models (LLMs) and large multimodal models (LMMs) offer transformative advantages in this regard. These models have demonstrated remarkable capabilities in generating human-like text, understanding complex queries, and providing contextually relevant responses. With interactivity, they can incorporate contextual inputs or user queries to provide tailored, situation-specific diagnoses. Furthermore, they can be designed to directly output decisions or actionable recommendations, eliminating the need for complex post-processing. More critically, LLMs shift the paradigm from low-information-density fault categories to high-information-density analysis, delivering detailed insights more than just defect categories, thereby enhancing both interpretability and practical utility in industrial applications.
Despite these advances, the application of large-scale models to fault diagnosis remains in its early stages, with several noteworthy contributions emerging recently. BearingFM [24] established a framework for training foundation models in this domain, utilizing domain knowledge-based data augmentation and contrastive learning to extract features from unlabeled vibration signals, achieving high accuracy with minimal labeled data. Tao et al. [25] leveraged the text modality by quantitatively selecting features of vibration signals to textualize the time series data and utilized Low-Rank Adaptation (LoRA) for fine-tuning LLMs, demonstrating how language models can interpret bearing conditions. Taking a different approach, FaultGPT [26] exploited the vision modality by extracting features from vibration time-frequency images, which were then paired with textual instructions to generate detailed diagnostic reports.
@@
Building upon these observations, we identify a critical oversight in current approaches: whether utilizing the text modality of LLMs or extending to the visual modality of LMMs, these methods fail to recognize a fundamental characteristic of bearing vibration signals, which is their intrinsic similarity to acoustic phenomena. Bearing vibration signals, as mechanical oscillations propagating through a medium, manifest as time-varying waveforms with frequency, amplitude, and temporal patterns that closely resemble those of sound waves. This acoustic likeness is evident in their shared spectral properties, such as harmonic peaks and resonance frequencies, which are routinely analyzed in audio processing but underutilized in fault diagnosis. Furthermore, despite the critical importance of aerospace propulsion systems and the severe consequences of their failure, large-scale models specifically designed for aero-engine bearing fault diagnosis have yet to be developed. The absence of such specialized models represents a significant gap in the current research landscape and limits the effectiveness of predictive maintenance in aerospace applications.
@@
Motivated by the above challenges, this paper aims to align general domain audio knowledge with aero-engine bearing vibration patterns, and for aerospace engine bearing fault diagnosis, proposes a novel framework, AeroGPT. As a large-scale audio model-based approach, AeroGPT addresses the limitations of current methods through two key innovations: Vibration Signal Alignment (VSA), which adapts general audio knowledge to the specific characteristics of aero-engine bearing vibrations, and Generative Fault Classification (GFC), which retains the generative capabilities of the underlying LLM to directly output interpretable fault labels, eliminating the need for post-processing. By leveraging the inherent generative capabilities of large-scale audio models, AeroGPT enables interactive, interpretable, and actionable fault diagnosis, greatly enhancing its practical utility in aerospace applications. Our experimental validation on an aerospace bearing dataset demonstrates that AeroGPT achieves an impressive 98.94% accuracy on the DIRG dataset and 100% classification accuracy on the HIT bearing dataset, outperforming traditional deep learning approaches. Additional qualitative analyses further validate the effectiveness of our approach to support interactive fault diagnosis and follow-up analysis, showcasing the interpretability and actionability of the generated outputs. These results underscore the potential of large-scale audio models to revolutionize fault diagnosis in critical aerospace systems.
@@
    To the best of our knowledge, this paper is the first to adapt large-scale audio model for industrial fault diagnosis, proposing a new paradigm that differs from the deep learning techniques previously used in this field.
@@
    This paper proposes Vibration Signal Alignment (VSA), which bridges the gap between general audio knowledge and domain-specific vibration patterns, enabling the model to adapt its acoustic representations to the specific characteristics of bearing vibration signal.
@@
    This paper proposes Generative Fault Classification (GFC), leveraging the inherent generative capabilities of LLMs to directly output interpretable, actionable insights rather than traditional logits that require post-processing.
@@
    This paper proposes AeroGPT, a novel framework based on large-scale audio model that transfers the knowledge in general audio domain to aero-engine bearing fault diagnosis while retaining the generative capabilities of the underlying LLM. Experimental results demonstrate that AeroGPT achieves state-of-the-art performance and highlight the potential of large-scale models to revolutionize fault diagnosis.
@@
The remainder of this paper is structured as follows: Section II introduces the methodology of the proposed AeroGPT framework, including the overview of the system architecture, audio knowledge acquisition process, LoRA-based domain knowledge adaptation techniques, Vibration Signal Alignment (VSA) mechanism, and Generative Fault Classification (GFC) approach; Section III presents the experimental validation, including comprehensive evaluations on the DIRG and HIT bearing datasets; Section IV concludes the paper by summarizing the contributions and findings.
II Methodology
II-A Overview of AeroGPT
@@
Figure 1: Demonstration of the overall framework of AeroGPT, along with its practical application in the aero-engine bearing fault diagnosis process.
@@
Figure 2: Technical components of the AeroGPT methodology. The framework is initialized with a foundation model, followed by Vibration Signal Alignment (VSA) to adapt general audio knowledge to domain-specific vibration patterns and Vibration Signal Alignment (GFC) to output interpretable fault labels.
@@
For aero-engine bearing fault diagnosis, this paper proposes AeroGPT, a novel framework leveraging large-scale audio models, and the overall pipeline is shown in Fig. 1. This framework explores for the first time the application of large audio models to aerospace bearing fault diagnosis, with the aim of generating interpretable and actionable diagnostic outputs directly from vibration signals. AeroGPT transforms the conventional classification-based fault diagnosis approach into an interactive, generative process capable of providing detailed insights beyond simple fault categorization.
@@
where $\mathcal{E}_{\text{audio}}$ and $\mathcal{E}_{\text{text}}$ are the audio and text encoders respectively, $\mathbf{v}_{\text{proc}}$ is the processed vibration signal, $\mathbf{p}$ is the prompt template, and $\mathbf{c}$ represents optional context information. The audio tokens and text tokens are then concatenated together as the input for the LLM, which then processes these tokens and generates a fault label as the output. Leveraging the generative capabilities of the LLM, analysis and follow-up questions can also be conducted interactively, allowing for a more comprehensive understanding of the fault condition.
@@
AeroGPT incorporates two key mechanisms that address fundamental challenges in utilizing large-scale audio models for fault diagnosis, as illustrated in Fig. 2. The first component, Vibration Signal Alignment (VSA), bridges the gap between general audio knowledge and the specific characteristics of aero-engine bearing vibrations. VSA processes paired inputs of vibration signals and corresponding textual descriptions, enabling the model to establish meaningful connections between acoustic patterns and their diagnostic interpretations. The second one, Generative Fault Classification (GFC), leverages the inherent capabilities of LLMs to directly generate diagnostic outputs in natural language rather than numerical logits that require post-processing. This approach provides actionable insights for maintenance personnel and supports interactive diagnostic sessions where follow-up questions can be addressed, demonstrating great practical utility of the system in aerospace applications. Detailed descriptions of the components are provided in the following sections.
II-B Audio Knowledge Acquisition
The acoustic understanding capabilities in this framework are developed through a comprehensive multi-stage process, as illustrated in the leftmost part of Fig. 2. This process systematically builds robust audio pattern recognition abilities that can be transferred to bearing vibration signal analysis.
The initial stage of knowledge acquisition involves exposure to diverse audio-related tasks. As shown in the diagram, this includes transcription tasks where audio inputs (represented by the microphone icon) are converted into textual outputs (e.g., “[en]Hello there!”). This process enables the model to develop fundamental representations of acoustic patterns and their corresponding semantic interpretations. By processing varied audio inputs through structured transcription tasks, the model learns to extract meaningful features from complex acoustic signals.
Building upon basic acoustic understanding, the second stage employs supervised fine-tuning (SFT) to enhance the model’s analytical capabilities. This stage focuses on higher-level audio understanding tasks, in which the model learns to address complex queries about audio content, such as “What does the speaker feel?” and “What is the sound?”, providing responses like “Very happy.”, “Sound of rain.” This stage cultivates the model’s ability to extract contextual information and identify subtle acoustic features. The model also learns to understand human intent through instruction pairs.

### v2

@@ Abstract
Aerospace engines, as critical components in aviation and aerospace industries, require continuous and accurate fault diagnosis to ensure operational safety and prevent catastrophic failures. While deep learning techniques have been extensively studied in this context, they typically output logits or confidence scores, necessitating post-processing to obtain actionable insights. Furthermore, the potential of large-scale audio models for this task remains largely untapped. To address these limitations, this paper proposes AeroGPT, a novel framework that transfers knowledge from the general audio domain to aero-engine bearing fault diagnosis. AeroGPT leverages a large-scale audio model and incorporates Vibration Signal Alignment (VSA) to adapt general audio knowledge to domain-specific vibration patterns, along with Generative Fault Classification (GFC) to directly generate interpretable fault labels. This approach eliminates the need for label post-processing and supports interactive, interpretable, and actionable fault diagnosis, thereby enhancing industrial applicability. Through comprehensive experimental validation on two aero-engine bearing datasets, AeroGPT achieves 98.94% accuracy on the DIRG dataset and 100% accuracy on the HIT bearing dataset, outperforming representative deep learning approaches. Qualitative analysis and further discussion also demonstrate its potential for interactive diagnosis and real-world deployment, highlighting the promise of large-scale audio models to advance fault diagnosis in aerospace applications.
@@ I Introduction
Aerospace engines serve as the cornerstone of modern aviation and aerospace industries, powering everything from commercial aircraft to space exploration vehicles [2, 23]. Their reliable operation is paramount to aviation safety, mission success, and the prevention of catastrophic failures that could result in significant human and economic losses. Within these sophisticated propulsion systems, aero-engine bearings are particularly critical components, operating under extreme conditions of high speeds, temperature variations, and significant mechanical stress [13, 24, 50]. The failure of these bearings can lead to severe engine malfunctions, making them a primary focus for condition monitoring and maintenance protocols [7, 23]. Fault diagnosis of aero-engine bearings has consequently emerged as an essential discipline, offering methodologies to detect, identify, and predict potential failures before they manifest as catastrophic events. Early and accurate bearing fault detection can significantly extend engine lifespan, optimize maintenance schedules, and ensure operational safety [42].
Traditionally, bearing fault diagnosis approaches rely on a two-step process of feature extraction followed by classification [30]. For instance, Medina et al. [28] proposed an approach using Poincare plots for feature extraction, with subsequent classification performed by a multi-class Support Vector Machine (SVM). However, such approaches are limited by their dependence on domain-specific feature engineering and conventional machine learning classifiers, which may fail to fully capture the complex dynamics of aero-engine bearing systems. While the subsequent rise of deep learning methods has addressed the challenge of manual feature engineering, they introduce a different limitation. These data-driven methods typically output logits or confidence scores, necessitating post-processing to derive actionable insights, whereas industrial applications require efficient and interpretable solutions.
@@
Recently, the emergence and development of large language models (LLMs) and large multimodal models (LMMs) offer transformative advantages in this regard. These models have demonstrated remarkable capabilities in generating human-like text, understanding complex queries, and providing contextually relevant responses. With interactivity, they can incorporate contextual inputs or user queries to provide tailored, situation-specific diagnoses. Furthermore, they can be designed to directly output decisions or actionable recommendations, eliminating the need for complex post-processing. More critically, LLMs shift the paradigm from low-information-density fault categories to high-information-density analysis, delivering detailed insights beyond defect categories, thereby enhancing both interpretability and practical utility in industrial applications.
@@
Building upon these observations, we identify a critical oversight in current approaches: whether utilizing the text modality of LLMs or extending to the visual modality of LMMs, these methods fail to recognize a fundamental characteristic of bearing vibration signals, which is their intrinsic similarity to acoustic phenomena. Bearing vibration signals, as mechanical oscillations propagating through a medium, manifest as time-varying waveforms with frequency, amplitude, and temporal patterns that closely resemble those of sound waves. This acoustic likeness is evident in their shared spectral properties, such as harmonic peaks and resonance frequencies, which are routinely analyzed in audio processing but underutilized in fault diagnosis. This observation reveals a significant methodological gap in current research: traditional deep learning methods implement a “signal $\rightarrow$ logits” pipeline, outputting probability distributions that require post-processing, while recent works leveraging large language models have adopted indirect approaches like converting vibration signals into time-frequency images for vision-language models, or textualizing signal features before feeding them to LLMs. However, no prior work has established a direct “signal $\rightarrow$ language” paradigm that preserves the native waveform structure. Furthermore, despite the critical importance of aerospace propulsion systems and the severe consequences of their failure, large-scale models specifically designed for aero-engine bearing fault diagnosis have yet to be developed. The absence of such specialized models represents a significant gap in the current research landscape and limits the effectiveness of predictive maintenance in aerospace applications.
@@
Motivated by the above challenges, this paper proposes AeroGPT, a novel framework that aligns general-domain audio knowledge with aero-engine bearing vibration patterns for fault diagnosis in aerospace engines. The core novelty of AeroGPT lies in modality-aligned transfer and task reformulation, which make large-scale audio modeling directly applicable to aero-engine bearing fault diagnosis by treating vibration signals as audio-like waveforms and leveraging the acoustic representation capability of audio foundation models. This design is motivated by the fact that bearing vibration signals are fundamentally mechanical oscillations with temporal–spectral characteristics that closely resemble acoustic phenomena. By exploiting this inherent similarity, AeroGPT can directly utilize rich pre-trained knowledge from audio foundation models without the information loss introduced by time–frequency image conversion or feature textualization. As a large-scale audio model-based approach, AeroGPT addresses the limitations of current methods through two key innovations: Vibration Signal Alignment (VSA), which adapts general audio knowledge to the specific characteristics of aero-engine bearing vibrations, and Generative Fault Classification (GFC), which preserves the generative capability of the underlying LLM to directly output interpretable fault labels, eliminating the need for post-processing. To enable efficient adaptation to the aero-engine bearing domain, we employ Low-Rank Adaptation (LoRA), a parameter-efficient fine-tuning method that freezes most pre-trained parameters and learns only low-rank updates, substantially reducing computational and memory costs while maintaining strong performance. By leveraging the generative capabilities of large-scale audio models, AeroGPT enables interactive, interpretable, and actionable fault diagnosis, enhancing its practical utility in aerospace applications. Experiments on two aero-engine bearing datasets demonstrate that AeroGPT achieves 98.94% accuracy on the DIRG dataset and 100% accuracy on the HIT bearing dataset, outperforming traditional deep learning approaches. Additional qualitative analysis and discussion further demonstrate the applicability and potential of AeroGPT to advance fault diagnosis in safety-critical aerospace systems.
@@
    To address the challenge that existing fault diagnosis methods rely on indirect signal representations and output non-interpretable logits, this paper is the first to adapt a large-scale audio model for industrial fault diagnosis, proposing a modality-aligned paradigm that directly processes vibration signals as audio-like waveforms and generates human-readable diagnostic outputs.
@@
    To bridge the domain gap between general audio knowledge and mechanical vibration patterns, this paper proposes Vibration Signal Alignment (VSA), an intermediate adaptation objective that explicitly learns vibration-language grounding through vibration-text pairs, enabling effective knowledge transfer to the aero-engine bearing domain.
@@
    To meet industrial requirements for interpretability and actionability, this paper proposes Generative Fault Classification (GFC), which leverages the inherent generative capabilities of LLMs to directly output structured, human-readable fault labels and support interactive follow-up analysis, eliminating the need for post-processing pipelines.
@@
    By integrating the above components, this paper proposes AeroGPT, a unified framework that achieves state-of-the-art performance on two aero-engine bearing datasets (98.94% on DIRG and 100% on HIT), demonstrating the effectiveness of the vibration-as-audio paradigm and the potential of large-scale audio models for aerospace fault diagnosis.
@@
The remainder of this paper is structured as follows: Section II reviews the related work on fault diagnosis; Section III details the methodology of the proposed AeroGPT framework, covering its system architecture, the audio knowledge acquisition process, LoRA-based domain knowledge adaptation, and the core mechanisms of Vibration Signal Alignment (VSA) and Generative Fault Classification (GFC); Section IV presents our experimental validation, including comprehensive evaluations on the DIRG and HIT bearing datasets; Section V provides further discussion on the computational considerations and safe integration strategies; finally, Section VI concludes the paper by summarizing our findings and outlining directions for future research.
II Related Work
The field of bearing fault diagnosis has evolved significantly, primarily driven by advances in machine learning techniques. This section reviews the related literature by categorizing prior work into two paradigms: the well-established discriminative approaches powered by deep learning, and the emerging generative approaches leveraging large-scale models.
II-A Discriminative Approaches with Deep Learning
The first paradigm, which has been the mainstream of research for the past few years, treats fault diagnosis as a discriminative classification task. The goal of these methods is to learn a mapping from raw signal inputs to a set of predefined fault labels by identifying discriminative features. This paradigm has been significantly empowered by data-driven methods, particularly deep learning approaches, which enable end-to-end learning directly from raw vibration signals [4, 46]. These methods automatically extract hierarchical features, eliminating the need for the manual feature engineering that traditional approaches require.
Convolutional Neural Networks (CNNs) have been widely adopted for this purpose, as they excel at capturing local patterns and spatial hierarchies in data. Wang et al. [40] proposed a one-dimensional CNN with an attention mechanism tailored for bearing fault enhancement and classification, establishing a new state-of-the-art performance on the wheelset bearing dataset. Apart from CNNs, Recurrent Neural Networks, especially Long Short-Term Memory (LSTM) networks, have also gained traction. LSTMs are adept at capturing temporal dependencies in sequential data, making them suitable for modeling the time-varying nature of vibration signals. Song et al. [36] developed a CNN-BiLSTM network that combines CNN’s spatial feature extraction capabilities with BiLSTM’s temporal modeling, demonstrating robust performance even with limited samples.
More recently, Transformer-based architectures [38], leveraging self-attention mechanisms to capture long-range dependencies and contextual information, have also emerged as powerful tools. For instance, Li et al. [19] proposed Dconformer, a novel CNN-Transformer network that combines joint-learning denoising with a multi-branch cross-cascaded architecture to extract both local and global features from noisy vibration signals. Xiang et al. [45] introduced a frequency channel-attention based Vision Transformer method, which integrates frequency domain characteristics with self-attention mechanisms. To address the high computational cost of Transformers, Guo et al. [9] developed SPCFormer, a lightweight Transformer variant with selective patches and channel modules designed for efficient fault diagnosis in resource-constrained systems like quadrotor helicopters.
Beyond these architectures, researchers have explored other neural paradigms to address specific industrial challenges. For instance, Autoencoders (AEs) are widely used for unsupervised feature learning [10, 8]. Guo et al. [8] developed a multivariate fusion covariance matrix network (MFCMN), which feeds specially constructed covariance matrices into a standard autoencoder to effectively handle multi-channel signals with limited labeled samples. In another innovative direction, Graph Neural Networks (GNNs) have been introduced to utilize inter-sensor dependencies [21, 20]. Li et al. [20] designed a spectral graph wavelet network (SGWN) capable of multiscale feature extraction while mitigating over-smoothing. Brain-inspired architectures like Spiking Neural Networks (SNNs) have also been explored, with Xu et al. [46] proposing a deep spiking residual shrinkage network (DSRSN) that achieves high accuracy and efficiency in noisy environments.
Alongside these architectural innovations, significant research has focused on learning strategies to enhance model adaptability and reduce data dependency. Prominent examples include transfer learning [22, 31], meta-learning [33], few-shot learning [39, 35], and unsupervised learning [43]. However, the aforementioned methods are fundamentally discriminative, outputting logits that require post-processing, which is a limitation that our generative approach aims to overcome.
II-B Generative Approaches with Large-Scale Models
A new paradigm is emerging with the advent of large-scale foundation models, which reformulates fault diagnosis as a generative task. Instead of merely classifying faults, these methods aim to generate rich, human-readable outputs, enabling interactive and interpretable diagnostics. The application of such generative models to industrial fault diagnosis remains in its early stages, but several noteworthy research directions are taking shape.
One major research thrust focuses on building comprehensive Industrial Foundation Models (IFMs) and enhancing their reasoning capabilities with structured knowledge. For instance, Ren et al. [34] proposed a system architecture for a general-purpose IFM, designed to handle diverse industrial modalities and tasks throughout a product’s lifecycle. To bolster the reliability of LLMs in complex industrial settings, other researchers have turned to knowledge graphs (KGs). Zhuang et al. [53] integrated a time-frequency KG with a large model to improve fault semantic capture from multimodal data. Similarly, Zhou et al. [52] developed CausalKGPT, a framework that enhances an LLM with a causal knowledge graph to perform cause analysis for quality problems in aerospace manufacturing. Nie et al. [29] further integrated LLMs with KGs and Retrieval-Augmented Generation (RAG) for fault reasoning and maintenance recommendations in CNC machine tools. These approaches aim to inject LLMs with deep, structured domain knowledge, making them more trustworthy and capable of complex reasoning.
A second, more task-specific direction involves developing multimodal diagnostic models that fuse different representations of machine health data. BearingFM [15] established a framework for training large-scale models in this domain, utilizing domain knowledge-based data augmentation and contrastive learning to extract features from unlabeled vibration signals, achieving high accuracy with minimal labeled data. Tao et al. [37] leveraged the text modality by quantitatively selecting features of vibration signals to textualize the time-series data and utilized LoRA for fine-tuning LLMs, demonstrating how language models can interpret bearing conditions. Taking a different approach, FaultGPT [1] exploited the vision modality by extracting features from vibration time-frequency images, which were then paired with textual instructions to generate detailed diagnostic reports. Wang et al. [41] proposed DiagLLM, which incorporates multimodal signal representations and expert knowledge to enable explainable bearing fault diagnosis. More recently, Li et al. [18] proposed FD-MVLLM, which combines raw time-series signals with their time-frequency image representations, reprogramming both modalities into a format digestible by an LLM to improve diagnostic accuracy. Zhang et al. [49] further explored discretized signal representations and lightweight adapters to bridge continuous vibration signals with the discrete token inputs of LLMs.
These pioneering works validate the potential of large-scale models in various industrial fields but also highlight the need for a more direct and modality-aligned approach. While IFMs and KG-enhanced models offer powerful reasoning, they cannot directly use the domain knowledge to perform signal-based fault diagnosis tasks. Meanwhile, existing multimodal diagnostic models typically rely on indirect signal representations like images, textualized features, or discretized tokens. Our work contributes to this emerging paradigm by being the first to treat vibration signals as audio, thereby addressing a key gap in the current state of the art.
III Methodology
III-A Overview of AeroGPT
@@
Figure 1: Overall framework of AeroGPT and its practical application to aero-engine bearing fault diagnosis.
@@
Figure 2: Technical components of the AeroGPT methodology. The framework is initialized with a foundation model, followed by Vibration Signal Alignment Stage to adapt general audio knowledge to domain-specific vibration patterns and Generative Fault Classification Stage to output interpretable fault labels.
@@
For aero-engine bearing fault diagnosis, this paper proposes AeroGPT, a novel framework leveraging large-scale audio models, and the overall pipeline is shown in Fig. 1. This framework explores the application of large-scale audio models to aerospace bearing fault diagnosis for the first time, with the aim of generating interpretable and actionable diagnostic outputs directly from vibration signals. AeroGPT transforms the conventional classification-based fault diagnosis approach into an interactive, generative process capable of providing detailed insights beyond simple fault categorization.
@@
where $\mathcal{E}_{\text{audio}}$ and $\mathcal{E}_{\text{text}}$ are the audio and text encoders respectively, $\mathbf{v}_{\text{proc}}$ is the processed vibration signal, $\mathbf{p}$ is the prompt template, and $\mathbf{c}$ represents optional context information. The audio tokens and text tokens are then concatenated together as input to the LLM, which then processes these tokens and generates a fault label as the output. Leveraging the generative capabilities of the LLM, analysis and follow-up questions can also be conducted interactively, allowing for a more comprehensive understanding of the fault condition.
@@
AeroGPT incorporates two key mechanisms that address fundamental challenges in utilizing large-scale audio models for fault diagnosis, as illustrated in Fig. 2. The first component, Vibration Signal Alignment (VSA), bridges the gap between general audio knowledge and the specific characteristics of aero-engine bearing vibrations. VSA processes paired inputs of vibration signals and corresponding textual descriptions, enabling the model to establish meaningful connections between acoustic patterns and their diagnostic interpretations. The second component, Generative Fault Classification (GFC), leverages the inherent capabilities of LLMs to directly generate diagnostic outputs in natural language rather than numerical logits that require post-processing. This approach provides actionable insights for maintenance personnel and supports interactive diagnostic sessions where follow-up questions can be addressed, demonstrating practical utility of the system in aerospace applications. Detailed descriptions of the components are provided in the following sections.
III-B Audio Knowledge Acquisition
The acoustic understanding capabilities in AeroGPT are developed through a multi-stage process, as illustrated in the leftmost part of Fig. 2. The initial stage involves exposure to diverse audio-related tasks such as transcription, enabling the model to develop fundamental representations of acoustic patterns and their semantic interpretations. Subsequently, supervised fine-tuning (SFT) enhances the model’s analytical capabilities through higher-level audio understanding tasks, cultivating the ability to extract contextual information and identify subtle acoustic features.